### v1

# Large Language Model for Vulnerability Detection and Repair: Literature Review and Roadmap

Abstract.
The significant advancements in Large Language Models (LLMs) have resulted in their widespread adoption across various tasks within Software Engineering (SE), including vulnerability detection and repair. Numerous recent studies have investigated the application of LLMs to enhance vulnerability detection and repair tasks. Despite the increasing research interest, there is currently no existing survey that focuses on the utilization of LLMs for vulnerability detection and repair. In this paper, we aim to bridge this gap by offering a systematic literature review of approaches aimed at improving vulnerability detection and repair through the utilization of LLMs. The review encompasses research work from leading SE, AI, and Security conferences and journals, covering 36 papers published at 21 distinct venues. By answering three key research questions, we aim to (1) summarize the LLMs employed in the relevant literature, (2) categorize various LLM adaptation techniques in vulnerability detection, and (3) classify various LLM adaptation techniques in vulnerability repair. Based on our findings, we have identified a series of challenges that still need to be tackled considering existing studies. Additionally, we have outlined a roadmap highlighting potential opportunities that we believe are pertinent and crucial for future research endeavors.
copyright: none
1. Introduction
A software vulnerability refers to a flaw or weakness in a software system that can be exploited by attackers. Recently, the number of software vulnerabilities has increased significantly (TARGETT, 2022), affecting numerous software systems. To mitigate these issues, researchers have proposed methods for automatic detection and repair of identified vulnerabilities. However, traditional techniques, such as rule-based detectors or program analysis-based repair tools, encounter challenges due to high false positive rates (Shahriar and Zulkernine, 2012) and their inability to work for diverse types of vulnerabilities (Zhou et al., 2024a), respectively.
Recently, Large Language Models (LLMs) pre-trained on large corpus have demonstrated remarkable effectiveness across various natural language and software engineering tasks (Hou et al., 2023). Given their recent success, researchers have proposed various LLM-based approaches to improve automated vulnerability detection and repair, demonstrating promising outcomes for both detection and repair tasks (Steenhoek et al., 2023; Zhou et al., 2024a). LLM-based approaches for vulnerability detection and repair are increasingly attracting attention due to their potential to automatically learn features from known vulnerabilities and find/fix unseen ones. Furthermore, LLMs have the potential to utilize rich knowledge acquired from large-scale pre-training to enhance vulnerability detection and repair.
Despite the increasing research interest in utilizing LLMs for vulnerability detection and repair, to the best of our knowledge, there has been no comprehensive literature review summarizing on the state-of-the-art approaches, current challenges, and future directions in this field. To effectively chart the most promising path forward for research on the utilization of LLM techniques in vulnerability detection and repair, we conducted a Systematic Literature Review (SLR) to bridge this gap, providing valuable insights to the community. In this paper, we collected 36 primary studies over the last 6 years (2018-2024). We then summarized the LLMs used by relevant papers and classified various techniques used to adapt LLMs. We also discussed the key challenges associated with using LLMs for vulnerability detection and repair and proposed multiple potential research directions.
Related Literature Reviews. Researchers have undertaken a series of research endeavors concerning Machine Learning (ML) for source code vulnerability detection or repair (Hou et al., 2023; Ghaffarian and Shahriari, 2017; Lin et al., 2020; Wu et al., 2022; Zhang et al., 2023a). Hou et al. (Hou et al., 2023) conducted a systematic literature review on LLMs for SE. However, due to the extensive range of SE tasks to summarize, their review did not categorize the detailed utilization of LLMs in vulnerability detection and repair. In contrast, our literature review is more focused on vulnerability detection and repair. Ghaffarian et al. (Ghaffarian and Shahriari, 2017) studied vulnerability analysis and discovery approaches published till 2016. Lin et al. (Lin et al., 2020) reviewed vulnerability detection approaches utilizing deep learning till 2020. Wu et al. (Wu et al., 2022) investigated vulnerability detection approaches published before May 2022. However, their focus was not on LLMs and primarily covered general ML models such as recurrent neural networks. In contrast, our literature review includes papers published until March 2024 and focuses on LLMs, encompassing 15 distinct LLMs utilized in 36 relevant studies. Recently, Zhang et al. (Zhang et al., 2023a) reviewed learning-based automated program repair, which covered vulnerability repair as part of its study scope. In Different from Zhang et al., our literature review is more focused on LLMs and vulnerability repair: we included more recent studies, offered a detailed categorization of LLM usages in vulnerability repair, and discussed specialized future directions for vulnerability repair with LLMs.
In general, this study makes the following contributions:
  • We present a systematic review of recent 36 primary studies focusing on the utilization of LLMs for vulnerability detection and repair.
  • We offer a comprehensive summary of the LLMs utilized in the relevant literature and categorize various techniques used to adapt LLMs to do these two tasks (vulnerability detection and repair).
  • We discuss the key challenges associated with using LLMs for vulnerability detection and repair, and propose multiple potential research directions.
2. Preliminaries
Figure 1. Study Identification and Selection Process
2.1. Vulnerability Detection/Repair Formulation
In this study, we focus on source code vulnerability detection and repair. Here, we present the task formulations of these two tasks.
Vulnerability Detection. Vulnerability detection is typically framed as a binary classification task: $X_{i}\to Y_{i}$. Specifically, given an input source code function $X_{i}$, a model predicts whether the input function is vulnerable ($Y_{i}=1$) or non-vulnerable ($Y_{i}=0$).
Vulnerability Repair. Vulnerability repair studies using LLMs frame the task as a sequence-to-sequence problem: $X_{i}\to Y_{i}$. Specifically, given a vulnerable code snippet $X_{i}$, an LLM model generates the corresponding repaired code $Y_{i}$.
2.2. Large Language Models (LLMs)
The term Large Language Model (LLM) was introduced to distinguish language models based on their parameter size, specifically referring to large-sized pre-trained language models (Zhao et al., 2023). However, the literature lacks a formal consensus on the minimum parameter scale for LLMs (Wang et al., 2023). In this paper, we adopt the LLM scope division and taxonomy introduced by Pan et al. (Pan et al., 2023) and categorize the mainstream LLMs into three groups according to their architectures: 1) encoder-only, 2) encoder-decoder, and 3) decoder-only LLMs. We will provide a brief introduction to some representative LLMs for each category due to the space limit.
Encoder-only LLMs. Encoder-only LLMs are a type of neural network architecture that utilizes only the encoder component of the Transformer model (Devlin et al., 2018). In the SE domain, examples of encoder-only LLMs include CodeBERT (Feng et al., 2020), GraphCodeBERT (Guo et al., 2020), CuBERT (Kanade et al., 2019), VulBERTa (Hanif and Maffeis, 2022), CCBERT (Zhou et al., 2023b), SOBERT (He et al., 2023), and BERTOverflow (Tabassum et al., 2020).
Encoder-decoder LLMs. Encoder-decoder LLMs integrate both the encoder and decoder modules of the Transformer model (Vaswani et al., 2017). The encoder processes the input sentence, while the decoder generates the target output text/code. Prominent examples of encoder-decoder LLMs include PLBART (Ahmad et al., 2021), T5 (Raffel et al., 2020), CodeT5 (Wang et al., 2021), UniXcoder (Guo et al., 2022), and NatGen (Chakraborty et al., 2022).
Decoder-only LLMs. Decoder-only LLMs exclusively utilize the decoder module of the Transformer model to generate the target output text/code. The GPT series, including GPT-2 (Radford et al., 2019), GPT-3 (Brown et al., 2020), GPT-3.5 (OpenAI, 2022), and GPT-4 (OpenAI, 2023), stand as prominent implementations of this model series. Additionally, in the SE domain, there are numerous decoder-only LLMs specialized for code as well. Examples include CodeGPT (Lu et al., 2021), Codex (Chen et al., 2021), Polycoder (Xu et al., 2022), Incoder (Fried et al., 2023), CodeGen series (Nijkamp et al., 2023c, a), Copilot (GitHub, 2023), Code Llama (Meta, 2023), and StarCoder (Li et al., 2023a).
3. Review Methodology
3.1. Research Question
In this paper, we focus on investigating the following three Research Questions (RQs):
  • RQ1: What LLMs have been utilized to solve vulnerability detection and repair tasks?
  • RQ2: How are LLMs adapted for vulnerability detection?
  • RQ3: How are LLMs adapted for vulnerability repair?
3.2. Search Strategy
As shown in Fig. 1, following the guide by Zhang et al. (Zhang et al., 2011), our initial step is to identify primary studies to answer the research questions (RQs) above. Because the first LLM (i.e., BERT (Devlin et al., 2018)) was introduced in 2018, our search focused on papers published from 2018 onwards (i.e., from January 1st, 2018, to March 1st, 2024). Next, we identified the top peer-reviewed and influential conference and journal venues in the domains of SE, AI, and Security. We included 13 conferences (ICSE, ESEC/FSE, ASE, ISSTA, CCS, S&P, USENIX Security, NDSS, AAAI, IJCAI, ICML, NIPS, ICLR) and 4 journals (TOSEM, TSE, TDSC, TIFS). After the manual searching, we identified 35 papers that were relevant to our research objectives.
In addition to manually searching primary studies from top-tier venues, we also conducted an automated search across 7 popular databases, including IEEE Xplore (IEE, [n. d.]), ACM Digital Library (ACM, [n. d.]), SpringerLink (Spr, [n. d.]), Wiely (Wie, [n. d.]), ScienceDirect (Sci, [n. d.]), Web of Science (Web, [n. d.]) and arXiv (arX, [n. d.]). The search string used in the automated search is crafted from the relevant papers identified in the manual search. Please kindly check the complete set of search keywords in our online appendix (app, 2024) due to the limited space. After conducting the automatic search, we collected 1,500 relevant studies with the automatic search from these 7 popular databases.
3.3. Study Selection
Inclusion and Exclusion Criteria. After the paper collection, we conducted a relevance assessment according to the following inclusion and exclusion criteria:
  • The paper must be written in English.
  • The paper must have an accessible full text.
  • The paper must be a peer-reviewed full research paper published either in a conference proceeding or a journal.
  • The paper must adopt LLM techniques to solve source code vulnerability detection or repair.
  • The paper has less than 5 pages.
  • Books, keynote records, panel summaries, technical reports, theses, tool demos papers, editorials, or venues not subject to a full peer-review process.
  • The paper is a literature review or survey.
  • Duplicate papers or similar studies authored by the same authors.
  • The paper does not utilize LLMs, e.g., using graph neural networks.
  • The paper mentions LLMs only in future work or discussions rather than using LLMs in the approach.
  • The paper does not involve source code vulnerability detection or repair tasks.
In the first phase, by filtering out short papers (exclusion criteria 1) and deduplication (exclusion criteria 4), the total number of included papers was reduced to 1,332. In the second phase, we manually examined the venue, title, and abstracts of the papers, and the total number of included papers declined to 212. Books, keynote records, panel summaries, technical reports, theses, tool demo papers, editorials, literature reviews, or survey papers were also discarded in this phase (exclusion criteria 2-3). In the third phase, we manually read the full text of the paper to remove irrelevant papers. Specifically, the vulnerability detection or repair papers, which do not utilize LLMs but other methods, e.g., graph neural networks (GNN) or recurrent neural networks (RNN), were dropped (exclusion criteria 5). We also excluded studies that do not focus on source code vulnerability, such as those on binary code, protocols, or network communication. Furthermore, we removed studies that just discussed LLM as an idea or future work (exclusion criteria 6). We also removed the papers focusing on other tasks rather than vulnerability detection or repair, such as vulnerable data generation, vulnerability assessment, etc. (exclusion criteria 7). After the third phase, we identify 61 primary studies directly relevant to our research topic.
Figure 2. Distribution of Publications per Year
Quality Assessment. To prevent biases introduced by low-quality studies, we formulated five Quality Assessment Criteria (see our online appendix (app, 2024)) to assess the clarity, validity, and significance of the papers. We employed a scoring system ranging from 0 to 3 (poor, fair, good, excellent) for each quality assessment criterion. Following the manual assignment of scores, we selected papers with total scores exceeding 12 (80% of the maximum possible score). After this quality assessment, we obtained 33 papers.
Forward and Backward Snowballing. To avoid omitting any possibly relevant work during our manual and automated search process, we also performed lightweight backward and forward snowballing  (Wohlin, 2014). This involved reviewing both the references cited in our selected 33 primary studies and the publications that cited these studies. As a supplement, we gathered 77 more papers and repeated the entire study selection process, including filtering, deduplication, and quality assessment, which resulted in the identification of three additional papers. Thus, we obtained a final set of 36 papers to study.
3.4. Data Extraction and Analysis
Fig. 2 illustrates the distribution of selected primary studies across each year. The earliest relevant study we identified was published in 2021 (Ziems and Wu, 2021). Subsequently, interest in exploring LLMs for vulnerability detection and repair has steadily increased, peaking in 2023, with 50% of the total studied papers. Additionally, it is noteworthy that in the first two months of 2024 alone, the number of published papers has already reached 10, constituting 27.8% of the total papers. This suggests a growing research interest in leveraging LLMs for vulnerability detection and repair. We also analyzed the publication venues of the included studies. ICSE emerges as the predominant conference venue for LLM studies on vulnerability detection and repair, contributing 27.8% of the total number of studies. Other notable venues are TSE (8.3%), FSE (5.6%), EMSE (5.6%), ISSTA (5.6%), SCAM (5.6%), and TOSEM (2.8%).
4. RQ1: What LLMs have been utilized?
Figure 3. Distribution of LLMs for Vulnerability Detection
Figure 4. Distribution of LLMs for Vulnerability Repair
In general, out of the 36 included studies, we identified 15 distinct LLMs that have been utilized.
LLMs Used for Vulnerability Detection. Fig. 3 illustrates the distribution of the LLMs utilized for vulnerability detection. CodeBERT (Feng et al., 2020) emerges as the predominant LLM in addressing vulnerability detection to date, representing 38.5% (15/39) of the use of LLMs in the included studies. Following closely, GPT-3.5 becomes the second most frequently studied model, accounting for 10.3% (4/39). Regarding the categories of LLMs, encoder-only LLMs comprise 61.5% (24/39) of the use of LLMs in the included studies. Following that, decoder-only LLMs account for 28.2% (11/39), and encoder-decoder LLMs constitute 10.3% (4/39).
LLMs Used for Vulnerability Repair. Fig. 4 illustrates the distribution of LLMs utilized for vulnerability repair. Unlike the detection task, CodeT5 (Wang et al., 2021) emerges as the predominant LLM in addressing vulnerability repair to date, representing 24.0% (6/25) of the use of LLMs in the included studies. Following closely, domain-specific pre-trained Seq2Seq Transformers become the second most frequently studied model, accounting for 16.0% (4/25) of the use of LLMs. Regarding the categories of LLMs, encoder-decoder LLMs comprise 48% (12/25) of LLMs used. Following that, decoder-only LLMs and encoder-only LLMs constitute 44% (11/25) and 8% (2/25) of the use of LLMs, respectively.
Answer to RQ1: Our analysis indicates that, to date, encoder-only LLMs have dominated vulnerability detection, while encoder-decoder LLMs have been prominent for vulnerability repair.
5. RQ2: How are LLMs adapted for vulnerability detection?
In RQ2, we shift our focus to examining the specific adaptation techniques of LLMs in vulnerability detection. Regarding the usage of LLMs, we summarize three major categories from the included studies: 1) fine-tuning (Devlin et al., 2018), which updates the parameters of LLMs using a labeled dataset, 2) zero-shot prompting (Brown et al., 2020), which freezes the model parameters and does not consider any labeled data, and 3) few-shot prompting (Brown et al., 2020), which also freezes the parameters but considers a few labeled example data points. As depicted in Fig. 5, 82% of the studies utilize fine-tuning, while 11% and 7% employ zero-shot prompting and few-shot prompting, respectively.
Figure 5. Adaptation Techniques of LLMs for Vul. Detection
In the following subsection, we will introduce detailed categories of LLM adaptation techniques for each of the following: 1) fine-tuning, 2) zero-shot prompting, and 3) few-shot prompting.
5.1. How are LLMs adapted for Vul. Detection?
5.1.1. Fine-tuning
The fine-tuning process usually involves several steps/stages, such as data preparation, model design, model training, and model evaluation. Regarding fine-tuning, we classify adaptation techniques into five groups based on the stages they mainly target: Data-centric innovations (data preparation), Combination with program analysis (data preparation), LLM+ other deep learning modules (model design), Domain-specific pre-training (model training), and Causal learning (training optimization).
Data-centric Innovations. Data-centric innovations focus on optimizing the vulnerability detection data used for fine-tuning LLMs. Prior studies (Yang et al., 2023; Croft et al., 2023; Nie et al., 2023; Chen et al., 2023a; Kuang et al., 2023) have found that existing vulnerability detection data can suffer from imbalanced label distribution, noisy or incorrect labels, and scarcity of labeled data. Researchers (Yang et al., 2023; Wen et al., 2023; Kuang et al., 2023) have explored how to address the data issues.
  • Data Sampling: To address label imbalance, i.e., having more non-vulnerable code samples than vulnerable ones in the dataset, Yang et al. (Yang et al., 2023) applied various data sampling techniques. They found that random oversampling on raw code data enhances the ability of the LLM-based vulnerability detection approach to learn real vulnerable patterns.
  • Positive and Unlabeled Learning: To address label quality issues such as noisy or incorrect labels, Wen et al. (Wen et al., 2023) proposed PILOT, which learns solely from positive (vulnerable) and unlabeled data for vulnerability detection. Specifically, PILOT generates pseudo-labels for selected unlabeled data and mitigates the data noise by using a mixed-supervision loss.
  • Counterfactual Training: To enhance the diversity of labeled data, Kuang et al. (Kuang et al., 2023) proposed perturbing user-defined identifiers in the source code while preserving the syntactic and semantic structure. This approach generates diverse counterfactual training data, which refers to hypothetical data (e.g., data after perturbing identifiers) differing from actual data (i.e., data without perturbation), useful for analyzing the effect of certain factors. Incorporating these counterfactual data enriches the training data for LLMs.
Combination of LLM with Program Analysis. Many LLMs undergo pre-training on extensive datasets through unsupervised objectives like masked language modeling (Devlin et al., 2018) or next token prediction (Radford et al., 2019). For those LLMs, they may prioritize capturing sequential features and could potentially overlook certain structural aspects crucial for understanding code. To address this limitation, several studies have proposed integrating program analysis techniques with LLMs. The idea involves utilizing program analysis to extract structural features/relations within code, which are then incorporated into LLMs to enhance their understanding. Specifically, Liu et al. (Liu et al., 2024) leveraged Joern (Yamaguchi, 2023) to build the AST and PDG of the function, leveraging this data to pre-train their LLM to predict statement-level control dependencies and token-level data dependencies within the function. Peng et al. (Peng et al., 2023) utilized program slicing to extract control and data dependency information, aiding LLMs in vulnerability detection. Wang et al. (Wang et al., 2024b) proposed to learn program presentations by feeding static source code information and dynamic program execution traces with LLMs. Additionally, Zhang et al. (Zhang et al., 2023c) proposed decomposing syntax-based Control Flow Graphs (CFGs) into multiple execution paths and feeding these paths to LLMs for vulnerability detection.
Combination of LLM with Other Deep Learning Modules. LLMs have their own inherent limitations. Firstly, most LLMs are based on the Transformer architecture, which primarily models sequential relations and features. Secondly, some LLMs (e.g., CodeBERT) impose restrictions on the length of input code snippets. For instance, the most frequently used LLM in vulnerability detection, CodeBERT, can only process 512 tokens. To address these limitations, researchers (Tang et al., 2023; Ziems and Wu, 2021) have attempted to combine LLMs with other DL modules:
  • LLM+GNN: To leverage the structural features of code more effectively, Tang et al. (Tang et al., 2023) introduced CSGVD, which employs graph neural networks (GNN) to extract the graph features of code and combine them with the features extracted by CodeBERT.
  • LLM+Bi-LSTM: To address the length constraints of LLMs, Ziems et al. (Ziems and Wu, 2021) first segmented the input code into multiple fixed-size segments. They then utilized BERT to encode each segment and incorporated a Bidirectional Long Short-Term Memory (Bi-LSTM) module to process the output of BERT on each segment. Finally, a softmax classifier was applied to the last hidden state of the Bi-LSTM to produce the final classification scores.
Domain-specific Pre-training. Domain-specific pre-training involves pre-training an LLM on data specific to a particular domain, such as vulnerability data, before fine-tuning it for a specific task within that domain. This process enables the LLM to better understand the data relevant to the domain. Several studies in the field of vulnerability detection have adopted this technique (Hanif and Maffeis, 2022; Liu et al., 2024; Ni et al., 2023; Wang et al., 2024b). These studies typically use one of three pre-training objectives:
  • Masked Language Modeling: This pre-training objective involves training the LLM to predict masked tokens in a corrupted code snippet. Hanif and Mahmood (Hanif and Maffeis, 2022) introduced VulBERTa, which pre-trained a RoBERTa (Liu et al., 2019) model on open-source C/C++ projects using the Masked Language Modeling objective. Since C/C++ is the programming language of the vulnerability detection data used to evaluate VulBERTa, this pre-training enhances VulBERTa’s ability to understand C/C++ code.
  • Contrastive Learning: This pre-training objective is to minimize the distance between similar functions while maximizing the distance between dissimilar functions. Through contrastive learning, the LLM can learn to capture the distinctive features of code. In particular, Ni et al. (Ni et al., 2023) and Wang et al. (Wang et al., 2024b) utilized different hidden dropout masks to transform the same input function into positive samples (i.e., those similar to the input function) while regarding other distinct functions as dissimilar samples. Then they pre-trained an LLM following the contrastive learning objective.
  • Predicting Program Dependencies: This pre-training objective aims to enhance the LLM by guiding the model to learn the knowledge necessary for analyzing dependencies in programs. Specifically, Liu et al. (Liu et al., 2024) pre-trained an encoder-only LLM on code from open-source C/C++ projects to predict the statement-level control dependency and token-level data dependency.
After the pre-training, those domain-specific pre-trained LLMs undergo fine-tuning on the vulnerability detection dataset to perform vulnerability detection.
Causal Learning. Although promising, LLMs have been found to lack robustness under perturbation or when encountering out-of-distribution (OOD) data (Mahbubur Rahman et al., 2024). Rahman et al. (Mahbubur Rahman et al., 2024) suggested that this weak robustness may be attributed to LLMs learning non-robust features, such as variable names, that have spurious correlations with labels. To address this issue, Rahman et al. proposed CausalVul, which first designs perturbations to identify spurious features and then applies causal learning algorithms, specifically do-calculus, on top of LLMs to promote causal-based prediction, enhancing the robustness of LLMs in vulnerability detection.
Figure 6. Effectiveness of Adaptation Techniques for Vulnerability Detection
5.1.2. Zero-shot Prompting
For zero-shot prompting, we only observe one adaptation technique (i.e., prompt engineering).
Prompt Engineering. Researchers (Fu et al., 2023; Zhou et al., 2024b; Purba et al., 2023) have attempted to devise effective prompts to guide LLMs in conducting vulnerability detection in the zero-shot prompting manner. Specifically, Zhou et al. (Zhou et al., 2024b) devised a prompt: “Now you need to identify whether a method contains a vulnerability or not. If has any potential vulnerability, output: ‘this code is vulnerable’. Otherwise, output: ‘this code is non-vulnerable’. The code is [code]. Let’s start:”. Fu et al. (Fu et al., 2023) proposed the prompt: “Predict Whether the C/C++ function below is vulnerable. Strictly return 1 for a vulnerable function and 0 for a non-vulnerable function without further explanation.” Purba et al. (Purba et al., 2023) devised a prompt: “[code] Is this code vulnerable? Answer in only Yes or No”. In these prompts, “[code]” refers to the input code.
5.1.3. Few-shot Prompting
For few-shot prompting, we only find one adaptation technique (i.e., retrieval augmentation).
Retrieval Augmentation. Retrieval augmentation is a technique used to enhance the few-shot prompting. It involves retrieving similar labeled data samples from the training set when given a test data sample and using these retrieved data as examples to guide the prediction of LLMs on the test sample. Liu et al. (Liu et al., 2023) proposed using efficient retrieval tools such as BM-25 and TF-IDF for retrieval. Zhou et al. (Zhou et al., 2024b) suggested using CodeBERT as a retrieval tool. This method first transforms code snippets into semantic vectors and then quantifies the similarity between two code snippets by calculating the Cosine similarity of their respective semantic vectors. It finally returned the top similar code based on the similarity scores.
5.2. How effective are the techniques used?
In this subsection, we examine the effectiveness of LLM adaptation techniques utilized in vulnerability detection. Specifically, we conduct a manual review of the ablation studies in the primary research to compare the performance of approaches with and without those LLM adaptation techniques. The differences observed can indicate the effectiveness of the LLM adaptation techniques. If a study does not provide such information, we exclude the results of the paper from this subsection.
Fig. 6 illustrates the effectiveness of LLM adaptation techniques in vulnerability detection. The blue dots represent the improvement rates reported in the primary studies, while the red triangle denotes the average improvement across the studies. When fine-tuning an LLM, domain-specific pre-training (DP), combination with program analysis (PA), combination with other DL modules (DL), data-centric innovations (DC), and causal learning (Causal) lead to improvements ranging from 3% to 16%, 4% to 16%, 1% to 8%, 2% to 17%, and 5%, respectively, in terms of Accuracy scores. For zero-shot prompting, we did not find a clear ablation study demonstrating the contribution of prompt engineering. For few-shot prompting, retrieval augmentation (RAG) has enhanced the effectiveness of LLMs, improving accuracy from 3% to 15%. These findings highlight the extent these methods were effective as reported in prior studies.
Answer to RQ2: Our analysis revealed three commonly used techniques to adapt LLMs for vulnerability detection: including fine-tuning ($\approx$82%), zero-shot prompting ($\approx$11%), and few-shot prompting ($\approx$7%). We noted that different adaptation techniques led to varying improvements, with combination with program analysis (PA) technique achieving the highest average improvement.
6. RQ3: How are LLMs adapted for vulnerability repair?
Regarding the usage of LLMs, we summarize three major categories from the included studies: 1) fine-tuning (approximately 80%), which updates the parameters of LLMs using a labeled dataset, 2) zero-shot prompting (approximately 13%), which freezes the parameters of LLMs and does not consider any labeled data, and 3) few-shot prompting (approximately 7%), which also freezes the parameters of LLMs but considers a few labeled examples.
In the following subsection, we will introduce detailed categories of LLM adaptation techniques for each of the following: 1) fine-tuning, 2) zero-shot prompting, and 3) few-shot prompting.
6.1. How are LLMs adapted for Vul. Repair?
6.1.1. Fine-tuning
The fine-tuning process usually involves several steps/stages, such as data preparation, model design, and model training. Regarding fine-tuning, we categorize adaptation techniques into four groups based on the stages they mainly target: Diverse relevant inputs (data preparation), Model-centric innovations (model design), Domain-specific pre-training (model training), and Reinforcement learning (training optimization).
Figure 7. Adaptation Techniques of LLMs for Vul. Repair
Diverse Relevant Inputs. Several studies have shown that in addition to the input vulnerable code, incorporating other diverse relevant inputs can boost the effectiveness of LLMs. Specifically, the AST of the vulnerable code input (Zhou et al., 2022), vulnerability descriptions (Zhou et al., 2022; Wei et al., 2023), and vulnerable code examples shared on CWE websites (Zhou et al., 2024a) can enhance the LLMs in repairing vulnerability. In addition, Wei et al. (Wei et al., 2023) found that vulnerability-inducing commits and vulnerability-fixing commits can also improve the effectiveness of the model. Lastly, Zhou et al. (Zhou et al., 2024a) found that certain vulnerable functions exceed the length limit of Transformer-based LLMs like CodeT5 (512 subtokens) (Zhou et al., 2024a). To address this limitation, they applied the Fusion-in-Decoder framework to partition a long code function into multiple segments and feed those segments into LLMs one by one. Their findings demonstrated that those segments can boost model effectiveness.
Model-Centric Innovations. Model-centric innovations encompass methodologies that prioritize revising the model architecture of LLMs (i.e., the Transformer (Vaswani et al., 2017)). Specifically, Fu et al. (Fu et al., 2024) drew inspiration from Vision Transformer (VIT)-based object detection techniques in computer vision. They proposed vulnerability queries within the pre-trained Transformer model to identify vulnerable code blocks within the source code. Additionally, they trained a model to learn a vulnerability mask, enhancing the attention of the vulnerability queries on the vulnerable code areas.
Domain-specific Pre-training. Domain-specific pre-training involves pre-training an LLM on data specific to a particular domain before fine-tuning it for the target task. Considering the similarity between bug-fixing and vulnerability-fixing tasks, several studies (Chen et al., 2023b; Chi et al., 2023; Zhang et al., 2023b; Zhou et al., 2024a) considered the task of bug-fixing as a pre-training task to enhance LLMs. Specifically, they first pre-trained LLMs on a bug fix corpus to fix bugs, and then fine-tuned LLMs on a vulnerability fix dataset to repair vulnerabilities. This kind of pre-training technique can be considered as Transfer Learning.
Reinforcement Learning. Islam et al. (Islam and Najafirad, 2024) introduced SecureCode, an LLM tuned using a reinforcement learning framework. This approach integrated syntactic and semantic rewards to generate fixes for vulnerable code. Specifically, they leveraged the CodeBLEU (Ren et al., 2020) score as the syntactic reward and BERTScore (Zhang et al., 2020) as the semantic reward. After combining these rewards, they applied the Proximal Policy Optimization (PPO) algorithm (Schulman et al., 2017) to fine-tune the CodeGen2-7B (Nijkamp et al., 2023b) model.
6.1.2. Zero-shot and Few-shot Prompting
For both zero-shot and few-shot prompting, we only observe one adaptation technique.
Prompt Engineering. For zero-shot prompting, Pearce et al. (Pearce et al., 2021) devised various prompts to guide LLMs in repairing vulnerabilities in a zero-shot manner. They incorporated error messages from static analysis tools like CodeQL or C sanitizers into their prompts, providing contextual information about the vulnerabilities. For instance, one prompt format is: ”// BUG: [vulnerability-description] //MESSAGE: [error-message] [Code] //FIXED VERSION:”, where [Code] represents the input vulnerable function. In addition, Wu et al. (Wu et al., 2023) also explored diverse prompts for directing LLMs in repairing vulnerabilities in a zero-shot manner. For example, they proposed a prompt that begins by commenting on the buggy code lines and adds ”BUG:” to highlight them. Then, the prompt appends ”FIXED:” at the end to guide the LLM in generating a corrected version of the code.
Furthermore, for few-shot prompting, Fu et al. (Fu et al., 2023) investigated a straightforward prompt strategy involving the provision of three repair examples within each prompt. This approach aimed to aid GPT-3.5 and GPT-4 in performing the repair task more effectively.
Figure 8. Effectiveness of Adaptation Techniques for Vulnerability Repair
6.2. How effective are the techniques used?
Fig. 8 illustrates the effectiveness of LLM adaptation techniques in vulnerability repair. The blue dots represent the improvement rates reported in the primary studies, while the red triangle denotes the average improvement across the studies. When fine-tuning an LLM, domain-specific pre-training (DP), combining LLM with diverse relevant inputs (DI), model-centric innovations (MC), and reinforcement learning (RL) have resulted in improvements ranging from 6% to 13%, 1% to 24%, 3% to 3.3%, and 6%, respectively, in terms of Accuracy scores. For zero-shot prompting and few-shot prompting, we did not find a clear ablation study demonstrating the contribution of prompt engineering. These results highlight the extent these methods were effective as reported in prior studies.
Answer to RQ3: Our analysis revealed three commonly used techniques to adapt LLMs for vulnerability repair: including fine-tuning ($\approx$80%), zero-shot prompting ($\approx$13%), and few-shot prompting ($\approx$7%). We found that different adaptation techniques resulted in varying degrees of improvement, with domain-specific pre-training (DP) achieving the highest average improvement.
7. The Road Ahead
In this section, we discuss current challenges (Section 7.1) and highlight promising opportunities (Section 7.2) for applying LLMs techniques in vulnerability detection and repair.
7.1. Challenges
Challenge 1: Lack of High-quality Vulnerability Dataset. One major challenge is the lack of high-quality vulnerability benchmarks. Previous studies (Croft et al., 2023; Nie et al., 2023; Chen et al., 2023a) have highlighted issues with existing vulnerability data, including noisy or incorrect labels (e.g., labeling clean code as vulnerable). This data quality issue is primarily attributed to the use of automatic vulnerability collection (Nie et al., 2023), which can gather large enough data for training DL-based models including LLMs but cannot ensure the complete correctness of the labels. While manually checking each data sample can ensure high quality, it is a very tedious and expensive process, especially when aiming for a large dataset. Constructing a high-quality vulnerability benchmark remains an open challenge to date. Moreover, there is a growing concern that the capabilities of LLMs may derive from the inclusion of evaluation datasets in the pre-training corpus of LLMs, a phenomenon known as data contamination (Jiang et al., 2024). To mitigate this concern, high-quality vulnerability benchmarks are preferred to have no overlap with the pre-training corpus of LLMs.
Challenge 2: Complexity in Vulnerability Data. Vulnerabilities can be inherently complicated, which brings challenges for their detection and repair with LLMs. For instance, inter-procedural vulnerabilities are prevalent in vulnerability data, and Li et al. (Li et al., 2024) discovered that detecting inter-procedural vulnerabilities poses greater challenges than intra-procedural ones. Moreover, vulnerabilities encompass a wide array of Common Weakness Enumeration (CWE) types (Zhou et al., 2023a), but LLMs may struggle with less frequent CWE types compared to frequent types (Zhou et al., 2024a, 2023a). In addition, vulnerabilities are typically represented in terms of code units, such as code lines, functions, or program slices within which the vulnerabilities occur. Sejfia et al. (Sejfia et al., 2024) observed a significant accuracy drop when detecting vulnerabilities that span multiple code units, such as spanning multiple functions. Future research should consider the complex nature of vulnerabilities when designing LLM-based solutions.
Challenge 3: Narrow Scope of Input Programs. Currently, LLM-based vulnerability detection and repair solutions primarily target the function level (or more fine-grained, at the line level). However, this narrow focus indicates that these approaches may not perform optimally when presented with a wider range of programs, such as classes or a whole repository. Function-level vulnerability detection approaches could overlook vulnerabilities that span multiple functions or classes (Sejfia et al., 2024). Similarly, function-level vulnerability repair approaches fall short when tasked with modifications across multiple functions within the repository (Zhou et al., 2024a). In addition to functions, future research could propose LLM-based detection/repair approaches capable of handling a broader range of programs.
Challenge 4: High Accuracy and Robustness. A vulnerability detection or repair solution with high accuracy is generally preferred, as it boosts developers’ confidence in the reliability of detections and fixes. However, current state-of-the-art approaches (Liu et al., 2024; Zhou et al., 2024a) have not yet achieved satisfactory accuracy, with 67.6% and 20% accuracy scores for vulnerability detection and repair, respectively. Moreover, the solution should maintain robustness against data perturbations or adversarial attacks to ensure its resilience. However, Yang et al. (Yang et al., 2022) and Rahman et al. (Mahbubur Rahman et al., 2024) discovered that LLMs are not robust against data perturbations. Future research should seek ways to improve the accuracy and robustness of LLM-based solutions.
Challenge 5: Establishing Trust and Synergy with Developers. To enhance developers’ trust, LLM-based approaches should explicitly notify developers when unable to provide accurate outputs, which allows developers to avoid spending time reviewing potentially inaccurate outputs (Le et al., 2015). In addition, there is limited interaction between developers and LLM-based vulnerability detection/repair solutions, which may hinder the establishment of trust and synergy during practical application. To bridge this gap, future research should investigate more effective strategies for fostering collaboration and trust between developers and LLM-based solutions (Lo, 2023). By nurturing trust and synergy, LLM-based solutions have the potential to evolve into intelligent allies, offering enhanced support to developers.
7.2. Opportunities
Opportunity 1: Curating High-quality Test Set. The absence of a high-quality vulnerability dataset poses a significant obstacle to vulnerability detection. While obtaining fully correct labels for a large dataset is expensive, a viable solution is to curate a high-quality test set (which is much smaller than the whole dataset) that can accurately assess progress in vulnerability detection. An easy approach for future work is to combine the manually checked vulnerability data samples scattered across several separate research works (e.g., (Croft et al., 2023; He and Vechev, 2023; Li et al., 2023b)) to form a high-quality test set. Future works can then do an empirical study to recognize the real progress in vulnerability detection with the high-quality test set. Furthermore, the community can maintain a living high-quality test set by adding new manually verified data to it. This curated test set can serve as a reliable benchmark for vulnerability detection.
Opportunity 2: Repo-level Vulnerability Detection/Repair. Current vulnerability detection and repair techniques primarily focus on the function or line level. One key reason is the input length limitations (512 subtokens) of the small LLMs like CodeBERT and CodeT5, which are predominantly used in existing studies. The 512 subtokens limitation aligns well with the function level data but faces difficulty in scaling up to classes or repositories. However, the advent of recent larger LLMs with significantly higher input length capacity, such as Code Llama (16,384 subtokens) and GPT-4 (24,576 sub-tokens), enables processing repo-level data more effectively. This presents an opportunity for future research to explore repo-level vulnerability detection/repair by leveraging these larger LLMs.
Opportunity 3: Using Larger Decoder-only LLMs. As presented in RQ1, it is evident that researchers favor encoder-only language models (61.5%) for vulnerability detection and encoder-decoder models (48.0%) for repair. Moreover, only 6 out of the 36 studies investigate LLMs with more than 7 billion parameters, suggesting a focus on smaller LLMs. This preference contrasts with a broader trend observed in LLMs for SE, as highlighted in a recent literature review (Hou et al., 2023). They found that since 2023, decoder-only models have become predominant, accounting for 73.1% of the papers studied in 2023. Additionally, among very large LLMs (e.g., ¿7B parameters), the majority are decoder-only models. Furthermore, decoder-only LLMs excel in generating code/text (OpenAI, 2023), making them well-suited for vulnerability repair, which is a generation task. While vulnerability detection is a classification task, it can be reframed as a generation task. For instance, decoder-only LLMs can be instructed to output ‘it is vulnerable’ given a vulnerable code and ‘it is non-vulnerable’ given a non-vulnerable code. Thus, large decoder-only LLMs can be leveraged to address both vulnerability detection and repair tasks. Currently, it seems that decoder-only LLMs, particularly those with parameters larger than 7 billion, are underutilized in vulnerability detection and repair, indicating untapped potential and offering opportunities for future work.
Opportunity 4: Advanced LLM Usage and Adaptation. Regarding LLMs usages, beyond the techniques observed in the included studies—such as fine-tuning and zero/few-shot prompting—there exist two more advanced usages of LLMs (Minaee et al., 2024) that have not been explored yet in vulnerability detection and repair: 1) LLM Agent: LLMs can serve as agents to decompose complex tasks into smaller components and employ multiple LLMs to address them (Wang et al., 2024a); 2) Usage of External Tools: LLMs can utilize external tools such as search engines, external databases, and other resources to enhance them (Qin et al., 2023).
Regarding detailed LLMs adaptation techniques, as illustrated in Fig.5 and Fig.7, the majority of proposed adaptations in this field are designed for fine-tuning. However, a plethora of advanced adaptation techniques for few-shot prompting remains unexplored. These techniques include iterative retrieval augmentation (Shao et al., 2023), recursive retrieval augmentation (Trivedi et al., 2022), and adaptive retrieval augmentation (Asai et al., 2023). Researchers can consider using those unexplored advanced LLM usages/adaptations in future works.
Opportunity 5: Customized LLMs for Vulnerability. Currently, widely used LLMs in vulnerability detection/repair are general-purpose LLMs (e.g., CodeBERT, CodeT5, and GPT-3.5) that do not fully exploit the wealth of open-source vulnerability data. A promising avenue is the development of customized LLMs tailored for vulnerability data. Some initial attempts in this direction include vulnGPT (Vicarius, 2023) and Microsoft Security Copilot (Microsoft, 2024). However, as these solutions are proprietary, their customized LLM details may not be fully disclosed. We advocate for collaborative efforts to develop open-sourced and effective customized LLMs for vulnerability.
8. Threats to Validity
The potential threat to validity is the risk of inadvertently excluding relevant studies during the literature search and selection phase. Incomplete summarization of keywords for vulnerability detection/repair or varied terminologies of LLMs may have caused relevant research studies to be missed in our review. To mitigate this risk, we initially performed a manual selection of 17 high-impact venues and extracted a relatively comprehensive set of standard keywords from relevant papers within these venues. In addition, we further augmented our search results by combining automated search with forward-backward snowballing.
9. Conclusion and Future Work
The use of Large Language Models (LLMs) for vulnerability detection and repair has been garnering increasing attention. This paper presents a systematic literature review of 36 primary studies on LLMs for vulnerability detection and repair, sourced from leading SE, AI, and Security conferences and journals. This review begins by analyzing the types of LLMs used in primary studies, shedding light on researchers’ preferences for different LLMs. Subsequently, we categorized a variety of techniques for adapting LLMs. Through our analysis, this review also identifies the challenges in this field and proposes a research roadmap outlining promising avenues for future exploration. In the future, we plan to broaden this literature review by incorporating additional vulnerability-related tasks, such as vulnerability localization and vulnerability assessment. Our online appendix is available at (app, 2024).